# Quasiparticle cooling algorithms for quantum many-body state preparation

### v1

@@ II.3 Rate equation
Note that the above derivation so far follows the steps of the textbook derivation of the Lindblad equation [50, 51], except here we focus on the Heisenberg-picture evolution of the occupation number operators, rather than on the system’s density matrix. However, to derive the Lindblad equation one typically assumes a rapid decay of bath temporal correlations, compared to system’s dynamical timescales; in contrast, in our analysis the memory effects of the bath will play an essential role.
To simplify the analysis, we make an assumption that the system’s density matrix in the beginning of a cooling cycle is diagonal in the quasiparticle basis, $[\hat{\rho}_{S},\hat{n}_{k}]=0$. Moreover, we will make a further assumption that the correlations between occupations of different levels can be neglected (closely related to molecular chaos assumption in kinetic theory). The former assumption can be justified in the limit $\theta\to 0$, by coarse-graining over sufficiently many cooling cycles (secular approximation). The latter assumption is more difficult to prove  [^2] We note that the second assumption is practically necessary, as the diagonal assumption still leaves us with $O(\exp(N_{S}))$ coupled equations to solve.. However, as we will show below, these assumptions lead to a theory of quasiparticle cooling which is in quantiative agreement with numerical simulations.
Under these two assumptions, the system’s density matrix is given by $\hat{\rho}_{S}=\sum_{\vec{\alpha}}\rho_{\vec{\alpha}\vec{\alpha}}\ket{\vec{\alpha}}\bra{\vec{\alpha}}$ |, where |$\ket{\vec{\alpha}}$ form a basis of states with occupation numbers $\{\alpha_{k}\}$, $\alpha_{k}=0,1$ and quasienergy
$\varepsilon(\vec{\alpha})=\sum_{k}\epsilon_{k}\alpha_{k},$
where $\epsilon_{k}$ is quasienergy of quasiparticle level $k$. The probability $\rho_{\vec{\alpha}\vec{\alpha}}$ can be expressed in terms of average occupation numbers $n_{q}={\rm Tr}(\hat{\rho}_{S}\hat{n}_{q})$ via
$\rho_{\vec{\alpha}\vec{\alpha}}=\prod_{q}n_{q}^{\alpha_{q}}(1-n_{q})^{1-\alpha_{q}}.$ (19)
Note that in writing this form, we used the assumption that there are no correlations between occupations of different quasiparticle states. This ansatz can also be viewed as a time-dependent Generalized Gibbs ensemble (GGE) ansatz for $\hat{\rho}$ [53, 54].
Next, we compute the change in the occupation number over one cycle, $\delta n_{k}$ using the above form of the density matrix and Eq. (16). Since density $\hat{n}_{k}$ commutes with $\hat{\rho}_{S}$, the Lamb shift contribution plays no role. Denoting the Bohr frequencies by
@@
$\delta n_{k}=\theta^{2}\sum_{\vec{\alpha},\vec{\beta}}\sum_{a,b,\tau_{1},\tau_{2}}\rho_{\vec{\alpha}\vec{\alpha}}(\beta_{k}-\alpha_{k})f_{\tau_{1}}f_{\tau_{2}}\Gamma_{ab}(\tau_{1}-\tau_{2})e^{i(\tau_{1}-\tau_{2})\Delta(\vec{\alpha},\vec{\beta})}\bra{\vec{\alpha}}\hat{A}^{a\dagger}\ket{\vec{\beta}}\bra{\vec{\beta}}\hat{A}^{b}\ket{\vec{\alpha}}$ (21)
This equation illustrates that the quasiparticle number evolution is controlled by (i) spectral correlations of the bath, given by $\Gamma_{ab}(\tau_{1}-\tau_{2})$, (ii) the choice of the time-dependent coupling, $f_{\tau}$, and (iii) the matrix elements of the operators $\hat{A}^{a}$ entering auxiliary-system coupling. With our above choice of bath unitary (7) and coupling (10), the bath spectral function reads
$\Gamma_{\mathcal{A}_{j}\mathcal{A}_{l}}(\tau)=\bra{\vec{0}}\hat{\sigma}^{-}_{\mathcal{A}_{j}}(\tau)\hat{\sigma}^{+}_{\mathcal{A}_{l}}(0)\ket{\vec{0}}=\delta_{jl}e^{-i\pi h\tau},$ (22)
@@
$\delta n_{k}=\theta^{2}\sum_{j=1}^{N_{A}}\sum_{\vec{\alpha}\vec{\beta}}\rho_{\vec{\alpha}\vec{\alpha}}(\beta_{k}-\alpha_{k})|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}\hat{\sigma}_{\mathcal{Q}_{j}}^{+}\ket{\vec{\alpha}}|^{2}.$ (23)
where
@@
We refer to $F_{h,T}(\epsilon)$ as the filter function associated with $f_{\tau}$, essentially being its Fourier transform. Eq. (23) is a general rate equation for quasiparticle level occupation numbers. Next, we discuss how the auxiliary quasienergy $\pi h$ and the coupling $f_{\tau}$ can be tailored to remove quasiparticles from the system, driving it towards the ground state.
II.4 Two cooling protocols
@@
In Eq. (23) above we expressed the change in the occupation numbers in terms of the transition rates between many-body states $|\vec{\alpha}\rangle$ and $|\vec{\beta}\rangle$ induced by the coupling to auxiliaries. It is natural to consider the prethermal limit where quasiparticle quasienergies are close to the energies of quasiparticles of the corresponding prethermal Hamiltonian [39]. For example, transverse-field Ising model with arbitrary integrability-breaking local terms is in the prethermal regime as long as $|J|,|g|\ll 1$ and a similar condition is satisfied by the additional, integrability-breaking terms. In the prethermal regime, generally $|\epsilon_{k}|\ll\pi$. We note that in this limit, the heating processes due to Floquet evolution, and the associated energy non-conservation, are suppressed exponentially in $1/|g|,1/|J|$.
The above rate equation (23) includes both heating ($\Delta({\vec{\alpha},\vec{\beta}})<0$) and cooling ($\Delta({\vec{\alpha},\vec{\beta}})>0$) processes. To ensure effective cooling, the filter function should be engineered to satisfy (for $\epsilon>0$):
$\bigg{|}\frac{F_{h,T}(+\epsilon)}{F_{h,T}(-\epsilon)}\bigg{|}\gg 1.$ (25)
The prethermalization assumption allows us to neglect processes where quasienergy changes by ±$\pm 2\pi,\pm 4\pi$ etc, since the corresponding matrix elements are strongly suppressed. Below, we will introduce two protocols that realise Eq. (25) to different degrees.
II.4.1 Step-wise cooling protocol (SCP)
@@
In the first strategy, one aims to resonantly remove single quasiparticles from the system, by approximately matching auxiliary excitation quasienergy $\pi h$ to a value $\epsilon$ within the quasiparticle band, $\Delta<\epsilon<\Delta+W$, where $\Delta$ is the single-particle quasienergy gap and $W$ is the bandwidth. The function $f_{\tau}=1/T$ (26) is kept constant, therefore giving a ‘step’ pulse for $1\leq\tau\leq T$. We will refer to this approach as the step-wise cooling protocol (SCP).
@@
We note that the SCP is the strategy used in Ref. [38]. The filter function corresponding to the SCP gives, up to a phase factor,
@@
which for $T\gg 1$ has a sharp peak at $\epsilon=\pi h$, with a width of order $\delta\epsilon=\frac{2\pi}{T}$. Thus, in the SCP, the processes of quasiparticle removal with quasienergies $\epsilon\approx\pi h$ are favored approximately according to
@@
The SCP is efficient in systems with a sufficiently large quasiparticle gap, as demonstrated in [38]. However, as the gap is decreased, in particular when a quantum critical point is approached, the heating processes become more important. Even in a gapped phase, the suppression of heating is only algebraic in quasiparticle energy, which leads to non-zero quasiparticle density in the steady state. Therefore, SCP does not allow to prepare a many-body ground state with high fidelity.
II.4.2 Modulated cooling protocol (MCP)

### v2

@@ II.3 Rate equation
Note that the above derivation so far follows the steps of the textbook derivation of the Lindblad equation [52, 53], except here we focus on the Heisenberg-picture evolution of the occupation number operators, rather than on the system’s density matrix. However, to derive the Lindblad equation one typically assumes a rapid decay of bath temporal correlations, compared to the system’s dynamical timescales; in contrast, in our analysis the memory effects of the bath will play an essential role.
To simplify the analysis, we make an assumption that the system’s density matrix at the beginning of a cooling cycle is diagonal in the quasiparticle basis, $[\hat{\rho}_{S},\hat{n}_{k}]=0$. We further assume that the correlations between occupations of different levels can be neglected. The former assumption can be justified in the limit $\theta\to 0$, by coarse-graining over sufficiently many cooling cycles (secular approximation). The latter assumption is more difficult to prove, but is closely related to the molecular chaos hypothesis in kinetic theory. In the context of integrable quantum systems it is often referred to as the time-dependent Generalized Gibbs ensemble (GGE) ansatz for $\hat{\rho}_{S}$ [54, 55, 56]. As we will show below, these assumptions lead to a theory of quasiparticle cooling which is in quantiative agreement with numerical simulations.
The diagonal assumption gives $\hat{\rho}_{S}=\sum_{\vec{\alpha}}\rho_{\vec{\alpha}\vec{\alpha}}\ket{\vec{\alpha}}\bra{\vec{\alpha}}$ |, where |$\ket{\vec{\alpha}}$ form a basis of states with occupation numbers $\{\alpha_{k}\}$, $\alpha_{k}=0,1$ and quasienergy
$\varepsilon(\vec{\alpha})=\sum_{k}\epsilon_{k}\alpha_{k},$ (19)
where $\epsilon_{k}$ is quasienergy of quasiparticle level $k$. The diagonal assumption still leaves us with $2^{N_{S}}$ coupled equations to solve. The second assumption, however, implies that the probabilities $\rho_{\vec{\alpha}\vec{\alpha}}$ can be expressed as a product over the single-quasiparticle distribution functions:
$\rho_{\vec{\alpha}\vec{\alpha}}=\prod_{q}\frac{e^{-\alpha_{q}\lambda_{q}}}{1+e^{-\lambda_{q}}},$ (20)
where the $\lambda_{q}$ serve to parameterise the single-particle distribution. Eq. (20) is the familiar form for a gas of non-interacting fermionic particles. Equivalently, the probabilities can be written in terms of the average quasiparticle occupation numbers $n_{q}={\rm Tr}(\hat{\rho}_{S}\hat{n}_{q})$,
$\rho_{\vec{\alpha}\vec{\alpha}}=\prod_{q}n_{q}^{\alpha_{q}}(1-n_{q})^{1-\alpha_{q}},$ (21)
by using the relation $n_{q}=e^{-\lambda_{q}}(1+e^{-\lambda_{q}})^{-1}$.
Next, we compute the change in the occupation number over one cycle, $\delta n_{k}$, using the above form of the density matrix and Eq. (16). Since density $\hat{n}_{k}$ commutes with $\hat{\rho}_{S}$, the Lamb shift contribution plays no role. Denoting the Bohr frequencies by
@@
$\delta n_{k}=\theta^{2}\sum_{\vec{\alpha},\vec{\beta}}\sum_{a,\tau_{1},\tau_{2}}\rho_{\vec{\alpha}\vec{\alpha}}(\beta_{k}-\alpha_{k})f_{\tau_{1}}f_{\tau_{2}}\Gamma_{aa}(\tau_{1}-\tau_{2})e^{i(\tau_{1}-\tau_{2})\Delta(\vec{\alpha},\vec{\beta})}|\bra{\vec{\beta}}\hat{A}^{a}\ket{\vec{\alpha}}|^{2}.$ (23)
With our above choice of bath unitary (7) and of the $\hat{B}$ operators (10), the bath spectral function reads
$\Gamma_{aa}(\tau)=e^{-i\pi h\tau},$ (24)
@@
$\delta n_{k}=\theta^{2}\sum_{a}\sum_{\vec{\alpha}\vec{\beta}}\rho_{\vec{\alpha}\vec{\alpha}}(\beta_{k}-\alpha_{k})|F_{h,T}(\Delta(\vec{\alpha},\vec{\beta}))|^{2}|\bra{\vec{\beta}}\hat{A}^{a}\ket{\vec{\alpha}}|^{2}.$ (25)
Here, $\rho_{\vec{\alpha}\vec{\alpha}}$ is defined via Eq. (21) to yield a closed set of equations for the $N_{S}$ quasiparticle occupations, and we defined
@@
We refer to $F_{h,T}(\epsilon)$ as the filter function associated with $f_{\tau}$. In short, it acts to filter out transitions occuring at unwanted frequencies, such as heating processes.
Eq. (25) is a general rate equation for quasiparticle level occupation numbers. It shows that the quasiparticle evolution is controlled by (i) the filter function, which encodes spectral features of the bath and time-dependent coupling, and (ii) the matrix elements of the operators $\hat{A}^{a}$ entering the system-auxiliary coupling. In order for the cooling protocol to remove quasiparticles from the system and drive into the many-body ground state, both of these features should be properly engineered, as we discuss next.
II.4 Choice of operators coupling to auxiliary qubits
First, we discuss the choice of the system cooling operators $\hat{A}^{a}$, entering the coupling Eq. (8). For simplicity, below we will consider 1d translationally invariant systems, with annihilation operator of a quasiparticle with quasimomentum $k$ given by
$\hat{Q}_{k}^{-}=\frac{1}{\sqrt{N_{S}}}\sum_{j=1}^{N_{S}}e^{ikj}\hat{Q}_{j}^{-},$ (27)
where $\hat{Q}_{j}^{-}$ is an operator centered at site $j$.
The auxiliary couplings $\hat{A}^{a}$ should be chosen such that the matrix elements for removing quasiparticle excitations in Eq. (25) are maximized. In principle, this can be accomplished by choosing the operators $\hat{A}^{a}$ to be quasiparticle annihilation operators $\hat{Q}^{-}_{j}$. Then, the matrix element appearing in Eq. (25) reads,
$|\bra{\vec{\beta}}\hat{Q}^{-}_{j}\ket{\vec{\alpha}}|^{2}=\frac{1}{N_{S}}\sum_{k}(1-\beta_{k})\alpha_{k}\prod_{q\neq k}\delta_{\alpha_{q}\beta_{q}},$ (28)
and the change of the quasiparticle occcupation numbers over one cooling period (assuming one auxiliary qubit per system site) is given by
$\delta n_{k}=-\theta^{2}|F_{h,T}(\epsilon_{k})|^{2}n_{k},$ (29)
i.e. the quasiparticle occupations decay exponentially with a rate set by the coupling strength and the filter function. The ground state is reached in time $O(\theta^{-2})$.
In practice, the precise form of the $\hat{Q}_{j}^{-}$ operators for a generic many-body system is often not known. These operators may have support on multiple sites, with size of the order of the correlation length. Furthermore, in phases with topological excitations, the operators may not even be local in the qubit basis. A clear example is the antiferromagnetic phase of the TFIM: while the excitations are local in a basis of non-interacting fermions, in the qubit basis, these excitations are domain walls, whose removal requires acting on a string of qubits (see Sec. III).
This complicated structure of the $\hat{Q}_{j}^{-}$ operators poses an obstacle to choosing the cooling operators to coincide with annihilation operators – current hardware constraints dictate that the $\hat{A}^{a}$ operators should be relatively simple operators with support on at most a few neighbouring qubits. Despite this, the principle of adiabatic continuity [46, 57] implies that in gapped systems in which the quasiparticles are not topological, a choice of $\hat{A}^{a}$ operators can be made which is simultaneously hardware-efficient and which cools to the many-body ground state.
To illustrate this, suppose that the Hamiltonian of a gapped system in a topologically trivial phase can be written as (extension to Floquet systems in the prethermal regime is straightforward):
$\hat{H}^{\prime}=\hat{H}_{0}+\lambda\hat{V},$ (30)
where $H_{0}$ is a simple solvable Hamiltonian where quasiparticles are created by local operators with support strictly on a few nearby sites, and $\hat{V}$ is a perturbation away from the simple point. Decomposing a Hamiltonian $\hat{H}^{\prime}$ into above form is not unique, and we choose one such representation which minimizes the perturbation strength $\lambda$. For the TFIM example we have $H_{0}=-g\sum_{i}\hat{Z}_{i}$ and $\hat{V}=\sum_{i}\hat{X}_{i}\hat{X}_{i+1}$, with $\lambda=J$; the quasiparticle annihilation operators of $\hat{H}_{0}$ are the single-spin operators $\hat{Q}_{0,j}^{-}=\hat{\sigma}^{+}_{j}$ in this case. According to Eq. (29), the non-interacting ground state at $\lambda=0$ can be prepared exactly by choosing the cooling operators $\hat{A}^{j}=\hat{Q}^{-}_{0,j}$, and the operators are strictly local in the qubit basis.
When the interactions are switched on ($\lambda\neq 0$), then as long as $\hat{H}^{\prime}$ and $\hat{H}_{0}$ remain in the same quantum phase, the principle of adiabatic continuity implies that the two Hamiltonians are related by a quasi-local unitary transformation [57]. Both the ground state and the quasiparticle operators become ‘dressed’ by the transformation, while retaining their quantum numbers. This dressing of quasiparticles by the interactions can be characterised by the spectral function
$S_{\hat{Q}^{-}_{0}}(k,\omega)=\sum_{\gamma^{\prime}}|\bra{\Omega^{\prime}}\hat{Q}_{0,k}^{-}\ket{\gamma^{\prime}}|^{2}\delta(\omega-(\varepsilon_{\gamma^{\prime}}-\varepsilon_{0})),$ (31)
where the sum runs over all excited states of the Hamiltonian $\hat{H}^{\prime}$ and $\Omega^{\prime}$ is now the interacting ground state with energy $\varepsilon_{0}$. In the $\lambda\to 0$ limit the spectral function exhibits a delta-function pole at the quasiparticle energy, $S(k,\omega)=\delta(\omega-\epsilon_{k})$, while for finite $\lambda$, as long as the quasiparticle lifetime remains long, the spectral function retains a sharp feature centred around $\epsilon_{k}$ and with a width proportional to the inverse quasiparticle lifetime  [^2] In practice it is the spectral function that is measured in experiment and the validity of the quasiparticle description requires the existence of a sharp pole.. The weight (quasiparticle residue) $Z_{k}$ of the pole is directly related to the matrix elements between the interacting ground state and single-quasiparticle excited states, $Z_{k}=|\bra{\Omega^{\prime}}\hat{Q}^{-}_{0,k}\ket{k^{\prime}}|^{2}$ [46]. Thus, as long as the quasiparticle picture holds, $Z_{k}$ remains non-zero and choosing the cooling operators according to $\hat{A}^{j}=\hat{Q}^{-}_{0,j}$ should lead to effective cooling of the low-energy excitations, albeit with cooling rate reduced by the factor of the quasiparticle residue.
This argument relies on the phase being adiabatically connected to a fixed-point Hamiltonian with simple, local quasiparticles, which is the case for gapped, topologically-trivial phases. In contrast, in systems where the underlying excitations are fractionalised, e.g. in the TFIM antiferromagnet or topologically ordered system, it is not possible to adiabatically connect the quasiparticle operators to local operators. Typically however, in such systems pairs or higher multiplets of quasiparticles may be annihilated locally. Similarly, as a system is tuned toward a critical point, the quasiparticle description may break down, leading to a vanishing quasiparticle residue, and a continuum of low-energy excitations emerges. In both of these cases we do not expect that the ground state can be prepared in constant time (scaling in the size of the system). We will see this explicitly in the case of the antiferromagnetic TFIM in Section III.
To summarise our discussion, a possible strategy for choosing local cooling operators is to adiabatically connect the system’s (effective) Hamiltonian to the nearest simple solvable Hamiltonian having local quasiparticle operators, and choose the coupling to auxiliary qubits to annihilate these quasiparticles. We note that, although above we viewed $\hat{V}$ as a perturbation, our argument is not restricted to weak-coupling: in Sec. IV, we will show that this method is effective in cooling an interacting spin ladder in the regime where the perturbation $\lambda$ becomes of order one.
II.5 Choice of the filter function and two cooling protocols
We now turn to the second essential ingredient of efficient cooling, the engineering of the filter function in Eq. (26). When the cooling operators are chosen according to the argument in the previous Subsection, the operators will still generally have small matrix elements for transitions that increase the system energy, i.e. those corresponding to $\Delta(\vec{\alpha},\vec{\beta})<0$ in Eq. (25). As we now discuss, a proper choice of the filter function corrects for this deficiency by suppressing negative frequency (heating) transitions.
The filter function can be seen as a Fourier summation of the time-varying function $f_{\tau}$, truncated to a finite number of terms, and so acts as a spectral filter on allowed transitions. Since in experiment the reset time is finite, spectral features of the filter function are broadened, with width of order $\delta\epsilon=1/T$; however, assuming that $T$ is sufficiently large compared to the inverse many-body gap, $T\Delta\gg 1$, it should be possible to choose the modulated function $f_{\tau}$ so that the filter function strongly suppresses heating processes,
[…]
@@ II.5 Choice of the filter function and two cooling protocols
For Floquet systems, the quasienergy is $2\pi$-periodic and this statement needs clarification. It is natural to consider the prethermal limit where quasiparticle quasienergies are close to the energies of quasiparticles of the corresponding prethermal Hamiltonian [39]. For example, the transverse-field Ising model with arbitrary integrability-breaking local terms is in the prethermal regime as long as $|J|,|g|\ll 1$ and a similar condition is satisfied by the additional, integrability-breaking terms. In the prethermal regime, generally $|\epsilon_{k}|\ll\pi$, and heating processes due to Floquet evolution, and the associated energy non-conservation, are suppressed exponentially in $1/|g|,1/|J|$. The prethermalization assumption then allows us to neglect processes in Eq. (25) where quasienergy changes by ±$\pm 2\pi,\pm 4\pi$ etc, since the corresponding matrix elements are strongly suppressed.
@@
Next, we introduce two protocols which realise Eq. (32) to different degrees. In the first strategy, one aims to resonantly remove single quasiparticles from the system, by approximately matching auxiliary excitation quasienergy $\pi h$ to a value $\epsilon$ within the quasiparticle band, $\Delta<\epsilon<\Delta+W$, where $\Delta$ is the single-particle quasienergy gap and $W$ is the bandwidth. The function $f_{\tau}$ can be chosen to have a sufficiently broad profile, with width of order $1/W$, so that the filter function is centered at $\pi h$, with width of order $W$. A simple example is to use a constant pulse:
II.5.1 Step-wise cooling protocol (SCP)
The auxiliary quasienergy is matched to the quasiparticle band, $\Delta<\pi h<\Delta+W$, and the function $f_{\tau}=1/T$ (33) is kept constant, therefore giving a ‘step’ pulse for $1\leq\tau\leq T$. We will refer to this approach as the step-wise cooling protocol (SCP).
@@ II.5.1 Step-wise cooling protocol (SCP)
The SCP was the strategy used in the experiment of Ref. [38]. The filter function corresponding to the SCP gives, up to a phase factor,
@@
which for $T\gg 1$ has a sharp peak at $\epsilon=\pi h$, with a width of order $\delta\epsilon=\frac{2\pi}{T}$. Thus, in the SCP, the processes of quasiparticle removal in Eq. (25) with quasienergies $\epsilon\approx\pi h$ are favored approximately according to
@@
The SCP is efficient in systems with a sufficiently large quasiparticle gap and small bandwidth, as demonstrated in [38]; in this case the resonance peak of the filter function can be made broad enough to cover the entire quasiparticle band, while only having small weight at negative frequencies, suppressing heating processes. However, as the gap is decreased, in particular when a quantum critical point is approached, the heating processes become more important. Even in a gapped phase, the suppression of heating is only algebraic in quasiparticle energy, which leads to non-zero quasiparticle density in the steady state. Therefore, SCP does not allow for high fidelity state preparation in most cases. A second disadvantage is the need to tune the auxiliary quasienergy $\pi h$ to overlap with the possibly unknown quasiparticle band.
Instead of transitions occurring only at the natural resonant frequencies of the auxiliary qubits, transitions that lead to cooling may be engineered by strongly modulating the coupling $f_{\tau}$ in the time domain. In the task of cooling the system toward thermodynamic equilibrium (in particular the ground state), it is natural to consider those functions which satisfy the detailed balance condition [53]:
$\bigg{|}\frac{F_{h,T}(+\epsilon)}{F_{h,T}(-\epsilon)}\bigg{|}=e^{\epsilon\beta/2},$ (36)
where $\beta$ is the inverse temperature of the target equilibrium. The detailed balance condition ensures that heating processes are thermally suppressed, with the relative ratio of cooling to heating rates matching that in thermal equilibrium. Note that the SCP explicitly violates this condition, and thus drives the system into a non-equilibrium state with small quasiparticle density.
In order to construct a filter function which satisfies the detailed balance condition, we start by observing that the ‘Fermi’ function, $1+\tanh(\epsilon\beta/4)$, satisfies Eq. (36). However, the filter function must satisfy periodicity in the quasienergy argument, and so a correct choice is given by the $2\pi$-periodic ‘symmetrised’ Fermi function
$F_{\beta}(\epsilon)=\frac{\pi}{2A}\sum_{m=-\infty}^{\infty}e^{im\pi}\tanh([\epsilon-m\pi]\beta/4),$ (37)
where $A$ is a normalisation constant, and the function is symmetric about $\epsilon=\pi/2$. For $|\epsilon|\ll\pi$, this function looks like the Fermi function with ‘thermal’ broadening $\beta$, and satisfies the detailed balance condition. By decomposing Eq. (37) as a Fourier series (see Appendix A), we are led to the following coupling time-dependence:
II.5.2 Modulated cooling protocol (MCP)